[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jul 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Jul 29 19:58:15 UTC 2017


HSU Shift Log for Sat Jul 29 2017 15:58:09 GMT-0400 (EDT)

Shift: H

Forecaster: D Zelinsky

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Unable to edit previously saved .spf files in NAWIPS.  
Appears to be a permissions problem. Talked to Matt Sardi about this.

AWIPS-2/NCP issues: None

Other Techincal issues: None

Miscellaneous issues: Requested GOES-16 Meso-1 from 15Z to 03Z between  
Irwin and Hilary.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170729/aeb413cc/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list